OpenStudy (anonymous):

What is the solution of 3 lx + 5l <6? A. x ≤ -3 and x ≥ -7 B. No solution C. x ≥ -3 and x ≤ -7 D. x < -3 or x > -7

OpenStudy (texaschic101):

3 |x + 5| < 6 divide by 3 |x + 5| < 6/3 | x + 5 | < 2 (x + 5) < 2 -- subtract 5 from both sides x < 2 - 5 x < -3 -(x + 5) < 2 -x - 5 < 2 -- add 5 to both sides -x < 2 + 5 -x < 7 -- multiply by -1 to make x positive, and change inequality sign x > -7 Answer D

OpenStudy (anonymous):

Thanks for the help, it was wrong tho.

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

what ? It was wrong .....how ?

OpenStudy (texaschic101):

@ganeshie8 ....can you look at this please

ganeshie8 (ganeshie8):

this is absolutely right ! :) http://www.wolframalpha.com/input/?i=3*%7Cx+%2B+5%7C+%3C6

ganeshie8 (ganeshie8):

unless your question itself has some formatting error when u copy pasted from u homework page @Awesome_Naya

OpenStudy (anonymous):

well when i sumbited it, i got a zero for that question

OpenStudy (anonymous):

no, its the right question

ganeshie8 (ganeshie8):

doesnt matter, D is the right answer for this question !

ganeshie8 (ganeshie8):

maybe, take a snapshot of the original question and attach...

OpenStudy (anonymous):

idk how, i'm on a computer

OpenStudy (texaschic101):

I thought I was going crazy there for a minute...thanks ganeshie :)

ganeshie8 (ganeshie8):

np :) wolfram is there for everything lol :P

OpenStudy (texaschic101):

is there supposed to be an equal sign in the problem ? That would make a difference

OpenStudy (texaschic101):

if there is an equal sign in the problem, then the answer would be A

OpenStudy (anonymous):

no, its suposed to be less than & eaqual to, but i couldnt find that

ganeshie8 (ganeshie8):

lol, there settles the suspense ;)

OpenStudy (texaschic101):

so the problem is less then or equal to 6 ?? If so, then the answer is A. I can't make the equal sign with the inequality sign either, so I would put : 3 |x + 5| <= 6 that equal sign can make the difference between getting it right and getting it wrong

ganeshie8 (ganeshie8):

\(\large 3|x + 5| \le 6 \)

OpenStudy (anonymous):

oh, so whats the difference ?

OpenStudy (anonymous):

how'd you do that @ganeshie8

ganeshie8 (ganeshie8):

\(\large 3|x + 5| \le 6 \) is different from \(\large 3|x + 5| \lt 6 \)

OpenStudy (texaschic101):

showing off ganeshie...lol....I can't make that sign

ganeshie8 (ganeshie8):

hahah yes :)

OpenStudy (texaschic101):

Naya..if there is an equal sign in the problem, there will be an equal sign in the answer

OpenStudy (anonymous):

\( \large 3|x + 5| \le 6 \(

ganeshie8 (ganeshie8):

try this again now

OpenStudy (anonymous):

\( \large 3|x + 5| \lt 6 \)

ganeshie8 (ganeshie8):

copy paste below thing exactly in ur next reply : ``` \( \large 3|x + 5| \le 6 \) ```

ganeshie8 (ganeshie8):

try again now :)

OpenStudy (anonymous):

\( \large 3|x + 5| \le 6 \)

OpenStudy (texaschic101):

Bingo :) you got it

OpenStudy (anonymous):

i dont see the difference

OpenStudy (texaschic101):

the equal sign makes a difference. If there is an equal sign in the problem, there will be an equal sign in the answer. If there is no equal sign in the problem, then there will be no equal sign in the answer Understand ?

OpenStudy (anonymous):

yeah, i have to go now

OpenStudy (texaschic101):

c ya :) thanks again ganeshie